OpenStudy (anonymous):

integral of t/(t^4+2)?

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits \frac{t}{t^4+2}dt\quad=\quad \int\limits \frac{1}{(t^2)^2+2}(t\;dt)\]We need to make the substitution, \(\Large\bf\sf u=t^2\)